# A matrix-free convex limiting framework for continuous Galerkin methods with nonlinear stabilization

### v1

@@ 3.2 Conditional IDP properties
Let $\bar{u}^{e}$ be the intermediate cell average defined by (20). Then the validity of (24) is guaranteed under a subcell CFL condition
@@
Adapting the proof technique employed in [27, Thm. 5.15], we will express $\bar{u}^{e}$ as a convex combination of states that are evolved using the one-dimensional LLF scheme and, therefore, possess the IDP property.
Splitting the set $\mathcal{N}_{e}$ into $\mathcal{N}_{e}^{\partial}=\{i\in\mathcal{N}_{e}\,:\,\mathbf{x}_{i}\in\partial K_{e}\}$ and $\mathcal{N}_{e}^{0}=\mathcal{N}_{e}\backslash\mathcal{N}_{e}^{\partial}$, we represent the old average $u^{e}$ defined by (21) as follows:
$u^{e}=\frac{1}{s}\left[\frac{m_{0}^{e}}{|K_{e}|}u_{0}^{e}+\sum_{i\in\mathcal{N}_{e}^{\partial}}\frac{m_{i}^{e}}{|K_{e}|}u_{i}\right].$ (28)
In the case $\mathcal{N}_{e}^{0}\neq\emptyset$, we use $s=1$, $u_{0}^{e}=\frac{1}{m_{0}^{e}}\sum_{i\in\mathcal{N}_{e}^{0}}m_{i}^{e}u_{i}$ and $m_{0}^{e}=\sum_{i\in\mathcal{N}_{e}^{0}}m_{i}^{e}$. If $K_{e}$ has no internal nodes, which is the case for $\mathbb{P}_{1}/\mathbb{Q}_{1}$ and triangular $\mathbb{P}_{2}$ elements, we use $s=2$, $u_{0}^{e}=u^{e}$ and $m_{0}^{e}=|K_{e}|$, as in [27, Thm. 5.15].
@@
$\bar{u}^{e}=\frac{1}{s}\left[\frac{m_{0}^{e}}{|K_{e}|}\bar{u}_{0}^{e}+\sum_{i\in\mathcal{N}_{e}^{\partial}}\frac{m_{i}^{e}}{|K_{e}|}\bar{u}_{i}^{e}\right],$ (29)
@@
$\bar{u}_{i}^{e}$ $=u_{i}-\frac{s\Delta t_{e}}{m_{i}^{e}}|\mathbf{c}_{i,e}|(\mathcal{F}(u_{i},u_{i};\mathbf{n}_{i,e})-\mathcal{F}(u_{0}^{e},u_{i};\mathbf{n}_{i,e})),\quad i\in\mathcal{N}_{e}^{\partial},$ (30)
$\bar{u}_{0}^{e}$ $=u_{0}^{e}-\frac{s\Delta t_{e}}{m_{0}^{e}}\sum_{i\in\mathcal{N}_{e}^{\partial}}|\mathbf{c}_{i,e}|(\mathcal{F}(u_{0}^{e},u_{i};\mathbf{n}_{i,e})-\mathcal{F}(u_{0}^{e},u_{0}^{e};\mathbf{n}_{i,e}))$ (31)
@@
Introducing the one-dimensional IDP bar states (cf. [11, 18])
@@
we express relations (30) and (31) as follows (cf. [11]):
$\bar{u}_{i}^{e}$ $=u_{i}+\frac{s\Delta t_{e}}{m_{i}^{e}}|\mathbf{c}_{i,e}|\lambda_{0i}^{e}(\bar{u}_{0i}^{e}-u_{i}),\quad i\in\mathcal{N}_{e}^{\partial},$ (32)
$\bar{u}_{0}^{e}$ $=u_{0}^{e}+\frac{s\Delta t_{e}}{m_{0}^{e}}\sum_{i\in\mathcal{N}_{e}^{\partial}}|\mathbf{c}_{i,e}|\lambda_{0i}^{e}(\bar{u}_{0i}^{e}-u_{0}^{e}).$ (33)
If the fake time step $\Delta t_{e}$ satisfies condition (27) with
@@
then $\bar{u}_{i}^{e},\ i\in\mathcal{N}_{e}^{\partial}$ and $\bar{u}_{0}^{e}$ are convex combinations of the IDP states that appear on the right-hand sides of (32) and (33), respectively. This proves the claimed IDP property of the convex combination (29). $\square$
@@
and assuming that $\mathcal{N}_{e}^{0}\neq\emptyset$ for simplicity, we arrive at (29) with $s=1$ and
@@
The CFL bounds for these LLF updates can again be derived as in [11, 27] using representations in terms of $\bar{u}_{0i}^{e}$ and $\bar{u}_{i,ee^{\prime}}^{e}$. The latter bar state is defined as in (25) but using the Riemann data $\hat{u}_{i,ee^{\prime}}$ in place of $u_{j}$. We leave the calculation of $\Delta t_{e}^{\max}$ as an exercise to the reader. The flux-limited version of our method will be IDP under a simpler and weaker CFL condition.
@@
We call $\Delta t_{e}$ a fake time step because, similarly to graph viscosity coefficients $d_{ij}$ of edge-based LLF schemes [11, 31, 42], the value of $\Delta t_{e}^{\max}$ is determined at the level of spatial semi-discretization (19) by the local maximum speed and the distance between the nodes of the Bernstein finite element.
@@
The intermediate cell averages $\bar{u}^{e}$ defined by (36) with $\Delta t_{e}\leq\Delta t_{e}^{\max}$ are IDP for $\Delta t_{e}^{\max}$ given by (37). We omit the formal proof because (36) has the same structure as the finite volume and DG schemes considered in [13, 26, 27]. Therefore, the IDP property of $\bar{u}^{e}$ is guaranteed under the same macrocell CFL condition, which imposes the simple bound (37) on the time step.
@@ 4 Monolithic convex limiting
We are now ready to present a constrained high-order extension of the low-order IDP schemes analyzed in Section 3. Adopting the general design principles of monolithic convex limiting [25, 33, 34] and adapting them to our element-based setting, we consider semi-discrete schemes of the form
@@
using a correction factor $\alpha_{ee^{\prime}}=\alpha_{e^{\prime}e}$ to be defined below. The resulting flux-limited version of formula (20) reads (cf. [26, 29])
@@
$\sum_{i\in\mathcal{N}_{e}}f_{i}^{e}=0.$ (44)
@@
The application of $\beta_{e}$ to the element contributions $f_{i}^{e},\ i\in\mathcal{N}_{e}$ can be interpreted as slope limiting of Barth–Jespersen type [5] in the scalar case and of Zhang–Shu type [63, 64, 65] in the case of a hyperbolic system.
@@
Its IDP property is guaranteed under the global CFL condition (23) with fake time steps $\Delta t_{e}=\Delta t_{e}^{\max}$ defined by (37) if the flux limiter is activated (otherwise at least for $\Delta t_{e}$ satisfying condition (27) of Lemma 1).
It remains to choose the algorithms for calculating $\alpha_{ee^{\prime}}$ and $\beta_{e}$. The choices that we make in this work are based on our experience with convex limiting of FCT and MCL type (see [27, Ch. 4 and 6], [8, 14, 25, 26, 40]).
@@ 4.1 Flux limiting
Since our formula (41) for the intermediate cell average $\bar{u}^{e}$ has the structure of a fully discrete finite volume scheme, flux limiting can be performed using the localized FCT algorithms presented in [13, 26, 29]. In general, the correction factor $\alpha_{ee^{\prime}}\in[0,1]$ must satisfy the symmetry condition
@@
with the fake time step $\Delta t_{e}=\Delta t_{e}^{\max}$, where $\Delta t_{e}^{\max}$ is given by (37).
@@
where [26, 28, 29],[27, Sec. 5.3.1.1]
@@
Guermond et al. [9, 13] present a general line search algorithm for finding $\alpha_{ee^{\prime}}$ that ensures positivity preservation for quasi-concave functions of the conserved variables. An explicit formula for such $\alpha_{ee^{\prime}}$ can be derived if the IDP constraints are linear or can be replaced with linear sufficient conditions, as shown in [25, 27, 45] for the Euler equations of gas dynamics. A very elegant pressure fix was derived by Abgrall et al. [3, 57] using the general framework of geometric quasi-linearization [59]. We use this approach to enforce positivity preservation for the pressure of the Euler system in Section 5.
@@
In contrast to the FCT-type convex limiting algorithms presented in [9, 13], we constrain the intermediate cell averages (41) of the semi-discrete scheme (38) instead of auxiliary states $u_{ee^{\prime}}^{\rm FCT}$ that depend on the real (global) time step $\Delta t$ rather than a fake (local) time step $\Delta t_{e}$. Therefore, our limiting strategy is monolithic in the sense that $\alpha_{ee^{\prime}}$ is independent of $\Delta t$.
@@ 4.2 Slope limiting
The intermediate nodal states $\bar{u}_{i}^{e}$ of formula (45) can be constrained using one of the element-based convex limiting algorithms reviewed in [27, Ch. 4 and 6]. The localized FCT schemes developed in [7, 8, 40] define
@@
To preserve global or local bounds $u_{i}^{\min}\in\mathcal{G}$ and $u_{i}^{\max}\in\mathcal{G}$ in the scalar case, we calculate $\beta_{i,e}$ using the FCT formula [8, 16, 29, 40]
@@
As noticed in [40, Sec. 4.3], this definition of $\beta_{i,e}$ has the same structure as the vertex-based version [24] of the Barth–Jespersen slope limiter [5].
In extensions to systems, the IDP property of $\bar{u}_{i}^{e}$ can be enforced as in the work of Zhang and Shu [63, 64, 65] or using explicit formulas derived from linear sufficient conditions. Examples of element-based algorithms using closed-form IDP limiters for the compressible Euler equations can be found in [8] and [27, Ex. 4.16 and 5.14]. The positivity fix presented in [3, 57] can also be adapted to the element-based format. The framework of geometric quasi-linearization [59] can be used to derive $\beta_{i,e}$ for general systems.
@@
The monolithic structure of the semi-discrete scheme (38) using the flux-corrected cell averages (41) is preserved by our slope limiting procedure. Similarly to the MCL schemes reviewed in [27], constraints are imposed at the level of spatial semi-discretization, which has a well-defined residual even in the steady state limit. This property distinguishes our approach from predictor-corrector FCT algorithms and classical slope limiters.
4.3 Matrix-free implementation
We have intentionally refrained from using matrix notation for our finite element schemes because all algorithmic steps can be performed without calculating global or local matrices. Even the system of equations (10) for the nodal time derivatives $\dot{u}_{j}$ can be solved using a matrix-free iterative method [2]. The residuals of fully discrete schemes can be assembled from element vectors without generating full element matrices. By contrast, high-order extensions of the FCT and MCL schemes proposed in [9, 25, 31] require costly matrix computations or subcell decomposition procedures that involve solution of local problems with sparse graph Laplacians [14, 33, 34, 40].

### v2

@@ 3.2 Analysis in the one-dimensional case
Let $\mathcal{G}$ be a convex invariant domain. Suppose that
$u_{i},u_{j}\in\mathcal{G}\quad\Rightarrow\quad\bar{u}_{ij}\in\mathcal{G}$ (27)
and the time step $\Delta t$ satisfies the CFL-like condition
$\frac{\Delta t}{m_{i}}\sum_{j\in\mathcal{J}_{i}\backslash\{i\}}2d_{ij}\leq 1.$
Then the forward Euler step (26) is IDP in the sense that
$u_{j}\in\mathcal{G}\ \forall j\in\mathcal{J}_{i}\quad\Rightarrow\quad u_{i}^{\rm SSP}\in\mathcal{G}.$
Proof:
Under the assumptions of the lemma, $u_{i}^{\rm SSP}$ is a convex combination of $u_{i}\in\mathcal{G}$ and $\bar{u}_{ij}\in\mathcal{G},\ j\in\mathcal{J}_{i}\backslash\{i\}$. It follows that $u_{i}^{\rm SSP}\in\mathcal{G}$. $\square$
The assumption (27) of the above lemma is met for $\bar{u}_{ij}$ of the form (24) if the local CFL condition (25) holds for the pseudo time step $\Delta t_{e}$.
3.3 Analysis in the multidimensional case
Guermond and Popov [guermond2016] found that a representation in the form (26) is also possible for a multidimensional finite element counterpart of the local Lax–Friedrichs method. This low-order IDP scheme uses the coefficients $d_{ij}$ of a graph Laplacian operator for algebraic stabilization purposes. High-order finite element approximations can also be written in the form (26) with $\bar{u}_{ij}$ replaced by $\bar{u}_{ij}^{H}=\bar{u}_{ij}+f_{ij}/(2d_{ij})$. The raw antidiffusive fluxes $f_{ij}$ can then be limited in a way that ensures the IDP property. The monolithic convex limiting strategy introduced in [kuzmin2020] produces the best flux approximation $f_{ij}^{*}\approx f_{ij}$ such that $\bar{u}_{ij}^{*}=\bar{u}_{ij}+f_{ij}^{*}/(2d_{ij})\in\mathcal{G}$ whenever $u_{j}\in\mathcal{G}\ \forall j\in\mathcal{J}_{i}$.
In the aforementioned extensions of (26), the bar states $\bar{u}_{ij}$ and their flux-corrected counterparts $\bar{u}_{ij}^{*}$ are associated with pairs of neighbor nodes or cells. Multidimensional intermediate states $\bar{u}^{e}$ of our low-order method (22) are associated with elements and cannot be interpreted as averaged solutions of one-dimensional Riemann problems. Indeed, formula (20) resembles an explicit update for a cell average of a DG approximation. Such updates are genuinely multidimensional and more natural in the finite element setting. The representation of the SSP-RK stage (22) in terms of the intermediate cell averages $\bar{u}^{e}$ reveals structural similarity to positivity-preserving DG schemes [zhang2010b, zhang2011, zhang2012] and makes it possible to handle CG discretizations similarly.
[…]
(a) edge-based
(b) element-based
Figure 1: Intermediate states of convex decompositions on a triangular mesh.
It turns out that the IDP property of $\bar{u}^{e}$ is guaranteed if $\Delta t_{e}$ is sufficiently small. We state and prove this result in the following lemma.
Lemma 2
Let $\bar{u}^{e}$ be the intermediate cell average defined by (20). Then
$u_{i},u_{j}\in\mathcal{G}\quad\forall j\in\mathcal{N}_{e}\quad\Rightarrow\quad\bar{u}^{e}\in\mathcal{G}$ (28)
under a subcell CFL condition of the form
@@ 3.3 Analysis in the multidimensional case
Adapting the proof technique employed in [kuzmin2023, Thm. 5.15], we will express $\bar{u}^{e}$ as a convex combination of states that are evolved using the one-dimensional LLF scheme and possess the IDP property by Lemma 1.
To begin, we split the index set $\mathcal{N}_{e}$ into $\mathcal{N}_{e}^{\partial}=\{i\in\mathcal{N}_{e}\,:\,\mathbf{x}_{i}\in\partial K_{e}\}$ and $\mathcal{N}_{e}^{0}=\mathcal{N}_{e}\backslash\mathcal{N}_{e}^{\partial}$. In the case $\mathcal{N}_{e}^{0}\neq\emptyset$, we introduce the auxiliary state
$u_{0}^{e}=\frac{1}{m_{0}^{e}}\sum_{i\in\mathcal{N}_{e}^{0}}m_{i}^{e}u_{i},\qquad m_{0}^{e}=\sum_{i\in\mathcal{N}_{e}^{0}}m_{i}^{e}$ (30)
and represent the old average $u^{e}$ defined by (21) as follows:
$u^{e}=\frac{m_{0}^{e}}{|K_{e}|}u_{0}^{e}+\sum_{i\in\mathcal{N}_{e}^{\partial}}\frac{m_{i}^{e}}{|K_{e}|}u_{i}.$ (31)
@@
$\bar{u}^{e}=\frac{m_{0}^{e}}{|K_{e}|}\bar{u}_{0}^{e}+\sum_{i\in\mathcal{N}_{e}^{\partial}}\frac{m_{i}^{e}}{|K_{e}|}\bar{u}_{i}^{e},$ (32)
@@
$\bar{u}_{i}^{e}$ $=u_{i}-\frac{\Delta t_{e}}{m_{i}^{e}}|\mathbf{c}_{i,e}|(\mathcal{F}(u_{i},u_{i};\mathbf{n}_{i,e})-\mathcal{F}(u_{0}^{e},u_{i};\mathbf{n}_{i,e})),\quad i\in\mathcal{N}_{e}^{\partial},$ (33)
$\bar{u}_{0}^{e}$ $=u_{0}^{e}-\frac{\Delta t_{e}}{m_{0}^{e}}\sum_{i\in\mathcal{N}_{e}^{\partial}}|\mathbf{c}_{i,e}|(\mathcal{F}(u_{0}^{e},u_{i};\mathbf{n}_{i,e})-\mathcal{F}(u_{0}^{e},u_{0}^{e};\mathbf{n}_{i,e}))$ (34)
@@
Introducing the one-dimensional IDP bar states (cf. [guermond2016, hoff1979])
@@
we express relations (33) and (34) as follows (cf. [guermond2016]):
$\bar{u}_{i}^{e}$ $=\Big(1-\frac{\Delta t_{e}}{m_{i}^{e}}|\mathbf{c}_{i,e}|\lambda_{0i}^{e}\Big)u_{i}+\frac{\Delta t_{e}}{m_{i}^{e}}|\mathbf{c}_{i,e}|\lambda_{0i}^{e}\bar{u}_{0i}^{e},\quad i\in\mathcal{N}_{e}^{\partial},$ (35)
$\bar{u}_{0}^{e}$ $=\Big(1-\frac{\Delta t_{e}}{m_{0}^{e}}\sum_{i\in\mathcal{N}_{e}^{\partial}}|\mathbf{c}_{i,e}|\lambda_{0i}^{e}\Big)u_{0}^{e}+\frac{\Delta t_{e}}{m_{0}^{e}}\sum_{i\in\mathcal{N}_{e}^{\partial}}|\mathbf{c}_{i,e}|\lambda_{0i}^{e}\bar{u}_{0i}^{e}.$ (36)
Note that these updates are of the form (26) and Lemma 1 is applicable. Let
@@
where $s\geq 1$. If the upper bound $\Delta t_{e}^{\max}$ of condition (29) is given by (37) with $s=1$, then the states $\bar{u}_{i}^{e},\ i\in\mathcal{N}_{e}^{\partial}$ and $\bar{u}_{0}^{e}$ are convex combinations of the IDP states that appear on the right-hand sides of (35) and (36), respectively. This proves the claimed IDP property of the convex combination (32).
If $\mathcal{N}_{e}^{0}=\emptyset$, which is the case for $\mathbb{P}_{1}/\mathbb{Q}_{1}$ and triangular $\mathbb{P}_{2}$ elements, we use $u_{0}^{e}=u^{e}$ and the convex decompositions (cf. [kuzmin2023, Thm. 5.15])
$u^{e}=\frac{1}{2}\left[u_{0}^{e}+\sum_{i\in\mathcal{N}_{e}}\frac{m_{i}^{e}}{|K_{e}|}u_{i}\right],\qquad\bar{u}^{e}=\frac{1}{2}\left[\bar{u}_{0}^{e}+\sum_{i\in\mathcal{N}_{e}}\frac{m_{i}^{e}}{|K_{e}|}\bar{u}_{i}^{e}\right].$
The states $\bar{u}_{i}^{e}$ and $\bar{u}_{0}^{e}$ are given by (35) and (36) with $\Delta t_{e}$ replaced by $2\Delta t_{e}$. It follows that the IDP property of $\bar{u}^{e}$ is guaranteed for $\Delta t_{e}\leq\Delta t_{e}^{\max}$, where $\Delta t_{e}^{\max}$ is the subcell CFL bound given by (37) with $s=2$. $\square$
@@
and assuming that $\mathcal{N}_{e}^{0}\neq\emptyset$ for simplicity, we arrive at (32) with
@@
The CFL bounds for these LLF updates can again be derived as in [guermond2016, kuzmin2023] using representations in terms of $\bar{u}_{0i}^{e}$ and $\bar{u}_{i,ee^{\prime}}^{e}$ with $e^{\prime}>E_{h}$. The latter bar state is defined as in (24) but using the Riemann data $\hat{u}_{i,ee^{\prime}}$ in place of $u_{j}$. We leave the calculation of $\Delta t_{e}^{\max}$ as an exercise to the reader. The flux-limited version of our method will be IDP under a simpler and weaker CFL condition.
@@
We call $\Delta t_{e}$ a pseudo time step because, similarly to graph viscosity coefficients $d_{ij}$ of edge-based LLF schemes [guermond2016, kuzmin2010a, luo1994], the value of $\Delta t_{e}^{\max}$ is determined at the level of spatial semi-discretization (19) by the local maximum speed and the distance between the nodes of the Bernstein finite element.
@@
The intermediate cell averages $\bar{u}^{e}$ defined by (39) with $\Delta t_{e}\leq\Delta t_{e}^{\max}$ are IDP for $\Delta t_{e}^{\max}$ given by (40). We omit the formal proof because (39) has the same structure as the finite volume and DG schemes considered in [guermond2019, kuzmin2021, kuzmin2023]. Therefore, the IDP property of $\bar{u}^{e}$ is guaranteed under the same macrocell CFL condition, which imposes the simple bound (40) on the time step.
@@ 4 Monolithic convex limiting
We are now ready to present a constrained high-order extension of the low-order IDP schemes analyzed in Section 3. Adopting the general design principles of monolithic convex limiting [kuzmin2020, kuzmin2020f, kuzmin2020a] and adapting them to our element-based setting, we consider semi-discrete schemes of the form
@@
using a correction factor $\alpha_{ee^{\prime}}=\alpha_{e^{\prime}e}$ to be defined below. The resulting flux-limited version of formula (20) reads (cf. [kuzmin2021, kuzmin2025a])
@@
$\sum_{i\in\mathcal{N}_{e}}f_{i}^{e}=0$ (47)
because
$m_{i}^{e}=\int_{K_{e}}\varphi_{i}\mathrm{d}\mathbf{x},\quad\sum_{i\in\mathcal{N}_{e}}\varphi_{i}\equiv 1,\quad\sum_{i\in\mathcal{N}_{e}}\nabla\varphi_{i}\equiv\mathbf{0},\qquad\sum_{i\in\mathcal{N}_{e}}\dot{u}_{i}\varphi_{i}=\dot{u}_{h}|_{K_{e}},$
$\sum_{i\in\mathcal{N}_{e}}\frac{m_{i}^{e}}{|K_{e}|}=1,\qquad\sum_{i\in\mathcal{N}_{e}}m_{i}^{e}u_{i}=\int_{K_{e}}u_{h}\mathrm{d}\mathbf{x}=|K_{e}|u^{e}=\Big(\sum_{i\in\mathcal{N}_{e}}m_{i}^{e}\Big)u^{e}.$
In the following theorem, we formally prove that our high-order CG-WENO discretization (10) can be recovered by adding the antidiffusive element contributions $f_{i}^{e}$ to the low-order scheme with deactivated flux limiter.
@@
The application of $\beta_{e}$ to the element contributions $f_{i}^{e},\ i\in\mathcal{N}_{e}$ can be interpreted as slope limiting of Barth–Jespersen type [barth1989] in the scalar case and of Zhang–Shu type [zhang2010b, zhang2011, zhang2012] in the case of a hyperbolic system.
@@
Its IDP property is guaranteed under the global CFL condition (23) with pseudo time steps $\Delta t_{e}=\Delta t_{e}^{\max}$ defined by (40) if the flux limiter is activated (otherwise at least for $\Delta t_{e}$ satisfying condition (29) of Lemma 2).
It remains to choose the algorithms for calculating $\alpha_{ee^{\prime}}$ and $\beta_{e}$. The choices that we make in this work are based on our experience with convex limiting of FCT and MCL type (see [kuzmin2023, Ch. 4 and 6], [dobrev2018, hajduk2021, kuzmin2020, kuzmin2021, lohmann2017]).
@@ 4.1 Flux limiting
Since our formula (44) for the intermediate cell average $\bar{u}^{e}$ has the structure of a fully discrete finite volume scheme, flux limiting can be performed using the localized FCT algorithms presented in [guermond2019, kuzmin2021, kuzmin2025a]. In general, the correction factor $\alpha_{ee^{\prime}}\in[0,1]$ must satisfy the symmetry condition
@@
with the pseudo time step $\Delta t_{e}=\Delta t_{e}^{\max}$, where $\Delta t_{e}^{\max}$ is given by (40).
@@
where [kuzmin2021, kuzmin2020b, kuzmin2025a],[kuzmin2023, Sec. 5.3.1.1]
@@
Guermond et al. [guermond2018, guermond2019] present a general line search algorithm for finding $\alpha_{ee^{\prime}}$ that ensures positivity preservation for quasi-concave functions of the conserved variables. An explicit formula for such $\alpha_{ee^{\prime}}$ can be derived if the IDP constraints are linear or can be replaced with linear sufficient conditions, as shown in [kuzmin2020, kuzmin2023, moujaes2025] for the Euler equations of gas dynamics. A very elegant pressure fix was derived by Abgrall et al. [abgrall-arxiv, wissocq2025] using the general framework of geometric quasi-linearization [wu2023]. A similar formula for $\alpha_{ee^{\prime}}$ is derived in Appendix A using the linearization proposed in [kuzmin2020, kuzmin2020c].
@@
In contrast to the FCT-type convex limiting algorithms presented in [guermond2018, guermond2019], we constrain the intermediate cell averages (44) of the semi-discrete scheme (41) instead of auxiliary states $u_{ee^{\prime}}^{\rm FCT}$ that depend on the (global) real time step $\Delta t$ rather than a (local) pseudo time step $\Delta t_{e}$. Therefore, our limiting strategy is monolithic in the sense that $\alpha_{ee^{\prime}}$ is independent of $\Delta t$.
@@ 4.2 Slope limiting
The intermediate nodal states $\bar{u}_{i}^{e}$ of formula (48) can be constrained using one of the element-based convex limiting algorithms reviewed in [kuzmin2023, Ch. 4 and 6]. The localized FCT schemes developed in [cotter2016, dobrev2018, lohmann2017] define
@@
To preserve global or local bounds $u_{i}^{\min}\in\mathcal{G}$ and $u_{i}^{\max}\in\mathcal{G}$ in the scalar case, we calculate $\beta_{i,e}$ using the FCT formula [dobrev2018, hajduk2020, kuzmin2025a, lohmann2017]
@@
As noticed in [lohmann2017, Sec. 4.3], this definition of $\beta_{i,e}$ has the same structure as the vertex-based version [kuzmin2010] of the Barth–Jespersen slope limiter [barth1989].
In extensions to systems, the IDP property of $\bar{u}_{i}^{e}$ can be enforced as in the work of Zhang and Shu [zhang2010b, zhang2011, zhang2012] or using explicit formulas derived from linear sufficient conditions. Examples of element-based algorithms using closed-form IDP limiters for the compressible Euler equations can be found in [dobrev2018, kuzmin2020c] and [kuzmin2023, Example 4.16]. The simple pressure limiter from [kuzmin2023, kuzmin2020c] is presented in Appendix A. The positivity fix employed in [abgrall-arxiv, wissocq2025] can also be adapted to the element-based format. The framework of geometric quasi-linearization [wu2023] can be used to derive $\beta_{i,e}$ for general systems.
@@
The monolithic structure of the semi-discrete scheme (41) using the flux-corrected cell averages (44) is preserved by our slope limiting procedure. Similarly to the MCL schemes reviewed in [kuzmin2023], constraints are imposed at the level of spatial semi-discretization, which has a well-defined residual even in the steady state limit. This property distinguishes our approach from predictor-corrector FCT algorithms and classical slope limiters.
4.3 Summary and discussion
The algorithm that we use to construct and limit a high-order CG-WENO discretization in each SSP-RK stage involves the following steps:
  1. 1.
    In a loop over elements $K_{e}$,
    • (a)
      Perform WENO reconstruction and calculate $\gamma_{e}$ given by (7).
    • (b)
      Calculate and store the contributions $s_{h}^{e}(\varphi_{i},u_{h}),\ i\in\mathcal{N}_{e}$ of (6).
    • (c)
      Add all contributions of $K_{e}$ to the right-hand side or residual.
  2. 2.
    Solve the linear system of equations (10) for the nodal states of $\dot{u}_{h}$.
  3. 3.
    In a loop over facets $S_{ee^{\prime}}$, calculate $f_{ee^{\prime}}=f_{ee^{\prime}}^{H}$ if flux limiting is deactivated and $f_{ee^{\prime}}$ given by (52) otherwise (see Section 4.1).
  4. 4.
    In a loop over elements $K_{e}$,
    • (a)
      Calculate the intermediate cell average $\bar{u}^{e}=u^{e}-\frac{\Delta t_{e}}{|K_{e}|}\sum_{e^{\prime}\in\mathcal{B}_{e}}|S_{ee^{\prime}}|f_{ee^{\prime}}$ using a pseudo time step $\Delta t_{e}\leq\Delta t_{e}^{\max}$ with $\Delta t_{e}^{\max}$ given by (37) if the flux limiter is deactivated and by (40) otherwise.
    • (b)
      Calculate the element contributions $f_{i}^{e},\ i\in\mathcal{N}_{e}$ given by (45).
    • (c)
      Use a slope limiter (see Section 4.2) to calculate $\beta_{e}$ given by (54).
    • (d)
      Calculate the intermediate nodal states $\bar{u}_{i}^{e},\ i\in\mathcal{N}_{e}$ given by (48).
  5. 5.
    Perform the SSP update (49) using a time step $\Delta t$ that satisfies (23).
[…]
Remark 4
The virtual finite element method developed by Abgrall et al. [abgrall2026] uses monolithic convex limiting in a “PAMPA” algorithm [abgrall2025] that also represents a semi-discrete scheme in terms of nodal and average states. The employed flux limiter constrains the cell averages at the time level $t^{n+1}$ using the MCL framework introduced in [kuzmin2020, kuzmin2021] and the pressure fix presented in [abgrall-arxiv, wissocq2025]. The raw antidiffusive element contributions $f_{i}^{e}$ of the slope limiting step correspond to a residual distribution method. Our element-based MCL formulation uses a different convex decomposition of a different baseline scheme (CG-WENO). Slope limiting is performed w.r.t. different states (intermediate cell averages rather than low-order nodal states). While our approach is designed for arbitrary-order finite element approximations on simplex or tensor-product meshes, the PAMPA scheme proposed in [abgrall2026] is suitable for general polygonal meshes and up to third-order accurate. The triangular mesh version [abgrall2025] uses $\mathbb{P}_{2}$ elements enriched by cubic bubble functions.